OpenStudy (anonymous):

10.0ml of 5.0% NaCl solution. To this you add 5.00ml of a 2.0% NaCl solution. What is the molarity of NaCl in this final mixture?

OpenStudy (amistre64):

dunno what molarity is but the final solution is: 10(.05) + 5(.02) = 15(x) .5 + .1 = 15x .6 = 15x x = .6/15 = 6/150 = 2/50 = 1/25 = .04 x = .04; which is 4%

OpenStudy (amistre64):

whats molarity? it sounds toothy lol

OpenStudy (anonymous):

molarity is moles of solute dissolved per liter, it's been so long since i've looked at lab math I'm completely lost

OpenStudy (amistre64):

well, we get 15mL of a 4% table salt solution...if that helps :)

OpenStudy (anonymous):

definitely, i appreciate the assistance

OpenStudy (amistre64):

youre welcome :)